FEL Upgrade Project Weekly Brief
November 4-8, 2002
We had another very productive week. Topping the list of major achievements includes getting rf in both cryomodules, one to full power with the other close behind. We got the gun assembled. We also got one of the optical cavity assemblies completed and ready to go into the tunnel (wait'll you see it, it's quite a piece of work!). We also put power in the wiggler and everything worked. There was also substantial progress in many other areas as detailed below.

All effort was spent this week in trying to resolve a severe cash flow crunch that has developed because of a significant curtailment of support from the State of Virginia, and, as of yet, non-receipt of JTO, AF, and most Navy funds for FY03.

We have submitted documentation to the Navy showing details of our technical status, plans and schedule for completion of the commissioning, and labor loading projected through FY03.

Nimel Theodore and Carlos Hernandez presented papers at the 49th AVS Symposium in Denver on our ion implantation process used for preparing the gun electrodes. Nimel's contribution won second place in the student competition.

WBS 4 (Injector):
Assembly of the ceramic stack and electrodes was completed. The stack was moved to the FEL vault and installed onto the gun chamber. A leak was found where the door operating bellows welds to its terminating block. An in-place repair of the weld was attempted and failed. The flange assembly with the bellows will be removed and the repair attempted again.

Gun HVPS - The leak in the high voltage connector was repaired. The HVPS tank is being pressurized again to check for any additional leaks. Several mounting holes on the conditioning resistor mechanism were moved to allow assembly. Final mechanical adjustments are underway.

WBS 6 (RF):
Quarter HVPS - Making good progress with unit 4 with the addition of 2 technicians from the EES group. Planning to connect to 480 VAC late next week and start checkout. The klystron cart for unit 4 is to be ready next week also.

Buncher - The Buncher was warmed up without RF and the water loop operated. All water fittings have been removed in preparation for the Gun and beam line bake. One fitting interfered with the mounting bracket and had to be cut off. A different style fitting is on order that will not interfere.

Zone 2 - SRF testing is complete with this zone. It operates at its old Drive High values. The new cavity tuning offset values are within +/- 5° of the old ones.

Zone 4 - SRF commissioning is making good progress. Nearly all cavities have been operated up to 10 MV/m. The klystron cathode voltage was set too low (10.8 kV) and is being raised (11.6 kV) to allow the required RF power. This was done, but several problems developed and are being addressed now.

WBS 8 (Instrumentation):
Trim magnet software and associated controls screens have been completed for the 1F to 3F region. Epics communications for the 66 kW PS has been checked out successfully. The PSS certification for the remaining large DC power supplies was completed. DC power has been connected to the Wiggler and will be checked out during the lock up 11/8/02. The new 5000 gallon bag for storage of sulfur hexafloride gas has been installed replacing the 2500 gallon bag used for the old gun HVPS. Testing continues for the HVPS interlock/controller chassis. This will be installed next week. The beam diagnostics for the 2G spectrometer line are being finalized. A holder for the Si wafer (2G view screen) is being finalized with the shop. A design for synchrotron light monitor attenuators is also being done with the shop. We are building from stock Al parts rather than using a previous design that required extensive machining.

The diagnostics for the 2G spectrometer line are being assembled. The design to hold a Si wafer (view screen) is being finalized with the shop. Work progressed on the Wiggler viewers with fabrication, assembly and installation of the BV Interface boxes air will be connected early next week and the alignment will begin.

Drive Laser Pulse Controller (DLPC) - The VME version of the ConOptics Model 305 is built. It picture of can be seen here:
http://laser.jlab.org/systems/wbs08_Instrumentation/drivelaser/dlpc/pcb-6uCO305/2002-11-08-pcb-6UCO305A-1.jpg. This plug-in module will provide the function that the Model 305 chassis does and will be fully integrated into the ioc (iocfel10). A link to a photo of the Model 305 chassis is here: http://laser.jlab.org/systems/wbs08_Instrumentation/drivelaser/chassis-CO305/ConOptics_Model-305.jpg. We are waiting on some rare RF connectors(SSMC, ~3 weeks) to finish the job. The Model 305 is currently operational and will stay online until the new system is up and running.

Drive Laser Controls - Bringing all the functions of the acromag card online continues. This multifunction I/O card does three functions, namely: discrete input and output (working), DAC control (working) & ADC input (in progress). Additionally, work continues on the RF/picomotor controller board. Parts are on order and the CAD work in in progress.

Identifying location and terminations of the signal and HV cables for the BLM system has begun in the 0F to 3F region. System prints were submitted to EECAD for the magnet communication upgrade as well as the new magnet configuration in the FEL. We continue to work with EECAD to clear up the backlogged drawings in their queue.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)

  • Magnetic success! Magnetic tests on the stepper stand show core field flatness to 1.5 parts in 10,000 over the good field region, field integral flatness within 2 parts in 10,000. Field integral is (with one field clamp fully in and the other side fully out--it was just a shake down run), within 0.12 % of the required value, almost within David's specification. The focus of the measurement group is now on these four magnets for an intense production magnetic measurement cycle.
  • All the bottoms of the 6 remaining magnets met the Purcell gap sheet piece flatness specification and those parts are now painted and have coils installed at MEI.
  • At the machine shop, tooling ball hole and pin placement on two of the six tops is now complete with the remainder close at hand. MEI is preparing for Purcell gap glue down, painting and then final assembly to ship in two weeks.

  • Injector Dipoles (DU/DV)
  • Two GVs and one GU were installed to complete the Extraction Chicane and the "Y" chamber inserted and hooked up to complete the vacuum system in that region.

  • Arc 180 Degree Dipoles (GY)
    Wang NMR is working to complete one magnet in 3.5 weeks.
  • They have reworked the coil mold to accommodate the coil inner bend down radius blend and are finishing sealing the coil (at the leads) in the mold for potting mid next week. They will use an automatically controlled, distributed resistance heater method to heat the mold and coil to the 275 °F curing temperature.
  • With respect to the inability to insert GG coils in the lower core because of partial coverage by the outer end return leg pieces, they have run into an additional problem. They encountered frozen bolts in the outer return legs that prevented their removal. Wang NMR will drill out the heads of the bolts to free up the parts.
  • Wang NMR is setting up to do the first gluing of Purcell gap shims late next week.

  • Arc Bend, Reverse Bend Dipoles (GQ, GX)
  • Two GQs and two GXs were completed by Wang NMR and are now inbound on a truck on the other side of Virginia slated for unloading Monday morning.
  • Wang vacuum potted the first of the new series of four coils for the remaining two GX magnets.

  • Quadrupoles
    Trim Quad (QT)
  • No change.

  • Sextupole (SF)
  • Master Machine received the original SCs for disassembly and are in a position to add the expansion pieces (most are complete) they have been making to enlarge the magnets to their "Commissioning Sextupole" configuration in two weeks.

  • Octupole (OT)
  • We are working on the procurement package to get these 2 magnets made in pieces by local shops. We have the coils already.

  • Beam Line and Vacuum
    The last chamber is expected near the end of November.
  • Chicane chambers continue fit up, welding and machining
  • The girders around the wiggler region were installed
  • The magnet measurement crew continued installation of the correctors.
  • LCW piping fabrication is at 85%.
  • The Dump line is now well defined and is being drawn up to use the 3 existing QI quadruples and dump module originally used in the IR Demo.



    WBS 10 (Wiggler):
    Ran the wiggler with its power supplies. Power supplies, PSS, etc., all worked great.

    WBS 11 (Optics):
    Note: The attached photos show Chris Behre and Joe Gubeli installing the foundation plate into the vacuum vessel, and then, the next day, Chris Behre and Mike Bevins installing the bulk of the optical cavity internals in the vacuum vessel.


    Our efforts this week were devoted to activities that will result in the optical cavity hardware, especially the vacuum vessels, being moved into the FEL vault next week. The "top hat" for the output coupler vacuum vessel passed its leak checked. Alignment of the internal components for the high reflector is nearly complete. This afternoon we will couple it's completed top hat to the lower vessel and leak check it while we begin installing the hardware inside the output coupler vessel.

    Other activities accomplished this week are:

    We'd like to take the opportunity to recognize one of our contract engineers, Erich Feldl. Erich was a long-time employee of the Lab, and has been on our staff since the beginning of the upgrade effort. Erich will consult for us from time to time, so we will continue to benefit from the decades experience he brought to the team.

    Commissioning/Operations Activities:
    See individual activities by WBS, especially the rf operation of the cryomodules.